OpenStudy (anonymous):

anyone know what factors can be multipied to get -35 but added to gett -18

OpenStudy (anonymous):

7 x 5 = 35

OpenStudy (anonymous):

however when added , they are not equal to 18

OpenStudy (anonymous):

yeah thought of that one but it doenst work on the second level

OpenStudy (anonymous):

x * y = -35 x + y = -18 x = -18 - y Plug back in: (-18-y)*y = -35 -18 y + y^2 = -35 y^2 - 18y + 35 = 0 y = around 15.78 ish or y = about 2.13 ish Then plug back for x.

OpenStudy (anonymous):

wait, isn't it supposed to be xy = 35 and x + y = 18?

OpenStudy (anonymous):

that's what she told me

OpenStudy (anonymous):

:'( Read the question.

OpenStudy (anonymous):

I know I did, but she told me it was + 35 and + 18 in the group chat

OpenStudy (anonymous):

mh...bk what's the question @_@

OpenStudy (anonymous):

sorry i did type it wrong first but i did change it in the group chat and in here'

OpenStudy (anonymous):

What is it? +38 and + 15 or -35 and -15?

OpenStudy (anonymous):

so I guess it's - 35 and - 18 :)

OpenStudy (anonymous):

Here are the exact values: \[x = -9 - 2 \sqrt{29}\] \[y = -9 + 2 \sqrt{29}\]